OpenStudy (anonymous):

Every 1/2 inch is 8miles if there are 6 miles, how many inches are there?

OpenStudy (anonymous):

I really suck with fractions sorry :(

OpenStudy (blurbendy):

what is 6/8 ?

OpenStudy (anonymous):

huh?

OpenStudy (blurbendy):

6/8 = 3/4 = .75 .75 * 1/2 = your answer

OpenStudy (anonymous):

ohh haha jeopardy style? LOL

OpenStudy (blurbendy):

yes... i'm trebek

OpenStudy (anonymous):

haha okay thank you!

OpenStudy (blurbendy):

yw